OpenStudy (anonymous):

Solve absolute value equation | 14x-1 | =1

OpenStudy (abb0t):

Solve: (14x-1) = 1 and -(14x-1) = 1

OpenStudy (anonymous):

is this question multiple choice by any chance?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Hold om

OpenStudy (anonymous):

A x=2 0r x 0 B x= 0 or -1/2 C x=0 or 1/2 D x = 2 or x -1/2

OpenStudy (anonymous):

Ok lets see….its not A or D hmmm

OpenStudy (anonymous):

I can't give the %100 exact answer the best i can say is its clear its not A or D.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Is that what it ended up being?

OpenStudy (anonymous):

No that's what I'm guessing cause it hast to be a positive and a negative

OpenStudy (anonymous):

Ok,Well if you got any more math questions that you need help with I'm always here.